Here's a function: K(r)=10r+3. I need help finding the following:
K(5)=
K(-6)=
K(6.8)=

Answer by jim_thompson5910(35256)

Simply plug the various values in place of r to get

K(5)=10(5)+3 = 50+3 = 53

So K(5)=53


K(-6)=10(-6)+3=-60+3=-57

So k(-6)=-57


I'll let you do the last one.